# Quantitative Introspection in Language Models: Tracking Internal States Across Conversation

### 19 Mar 2026 v1

@@ A.1 Conversation generation
For each model under study, we generated 40 multi-turn conversations of 10 turns each (one turn == one user message plus one assistant response), yielding 400 observation points per experimental condition. Conversations were generated using the model under study as the assistant and Gemini-2.5-Flash (via the OpenRouter API) as a simulated user. The simulated user receives a system prompt that instructs it to behave as a natural human user discussing a specified topic: it must stay on topic, keep messages brief (1–4 sentences), ask follow-up questions and share details, but not ask about the assistant or mention the simulation. The assistant model receives a minimal system prompt: You are a helpful assistant. You give very brief responses to the user’s questions (max: 5 sentences). The assistant model uses temperature =0.8=0.8, top-$p=0.9$, max 256 tokens. The user simulator uses temperature =0.7=0.7, top-$p=0.95$, max 256 tokens.
@@ A.3 Self-report query structure
The self-report questions follow the template: Hey, could you rate how [concept] you are [feeling] right now, from 0 to 9? Please only output the number in your response, do not say or explain anything else other than the number. The specific phrasings are:
@@
  • Impulsivity: “…rate how impulsive are you feeling right now, from 0 to 9…”
Self-report responses are generated with max 8 new tokens, temperature =0.8=0.8, top-$p=1.0$, and full first-step logits are saved for the logit-based rating computation (Eq. 1). The logit-based rating applies softmax over the 10 options, and computes the probability-weighted expected value.
@@ B.1 Training procedure
Each concept probe is trained on N=2024N=20-24 neutral questions, shared across both poles. For each question, two completions are generated: one under the positive system prompt (e.g., “You are a happy assistant…”) and one under the negative system prompt (e.g., “You are a sad assistant…”). Completions use greedy decoding with a maximum of 64 new tokens. Hidden-state activations are extracted after each transformer block (excluding the embedding output), and representations are computed as the mean over all assistant-response tokens.
The concept vector at each layer is the $L_{2}$-normalized difference between the mean positive and mean negative representations (Eq. 3). The best layer is selected as the one maximizing Cohen’s $d$ on a set of held-out evaluation texts (20 positive and 20 negative texts), with the search restricted to the 20th–80th percentile of layers.
@@ B.2 Concept pair system prompts
Below we list the complete system prompts used to train each of the four concept probes. Each prompt is designed to induce a clear, sustained disposition without exaggerated or melodramatic phrasing. Note that prompts are not exactly equivalent across concepts and contain specific intructions that were finetuned to the LLaMA 3B model, to make it act as desired.
@@ B.3 Training questions
Each concept probe is trained on a small set of neutral prompts, with between 20 and 24 training prompts depending on the concept. These prompts span factual and everyday scenarios selected to avoid strongly priming the target emotive state (see released code for the full lists).
@@ Appendix C Recovering informative self-reports
Appendix D Temporal structure and self-steering robustness
Fig. 7 expands the temporal and causal analyses from Fig. 3 by showing how the logit-based self-report metric changes across turns for each self-steering alpha condition. The goal here is to establish that the temporal effects discussed in the main text are not artifacts of a single summary statistic.
Figure 7: Supplementary self-steering drift curves for Fig. 3. Panels A–D show turn-wise logit-based self-reports under five steering strengths for wellbeing, interest, focus, and impulsivity; colored curves denote steering alpha and shaded bands denote cluster-bootstrap 95% CIs. Temporal drift persists under steering, but its magnitude depends on alpha.
Figs. 7A–D show the full self-steering drift curves, one concept per panel and one curve per steering strength (red corresponds to a more positive alpha, blue corresponds to more negative alpha). Notice how drift is stronger for more positive alpha values in the case of wellbeing and impulsivity (Figs. 7A and 7D), while it is stronger for more negative alpha values in the case of interest and focus (Figs. 7B and 7C). Also notice how negative alpha curves have lower mean self-report values, while positive alpha curves have higher mean self-report values, for all four concepts. This reflects the fact that our self-steering experiments show causal dependence between internal representation and self-report in all cases analyzed here.
@@ Appendix F Probe validation and cross-model generalization
Fig. 9 supports the scale and cross-family claims in Fig. 5 by giving more detail on probe quality and sign validation. Fig. 9A shows that probe quality varies strongly across LLaMA 1B, 3B, and 8B, even under a fixed training procedure. For example, for wellbeing concept probes, the maximum Cohen’s $d$ across layers was 1.7 for the 1B model, while it reached 3.4 for the 8B model. Furthermore, even though probe training was finetuned to the 3B model, probe quality was not consistently highest for that model. For example, for the interest concept probes, the 3B model shows a maximum Cohen’s $d$ across layers of 1.7, while the 1B and 8B models show substantially higher values (3.4 and 2.1, respectively). Even though there was variation between these values across models and concepts, in all cases probe training successfully achieved significant separation between evaluation examples in the semantically expected direction, with all best-layer tests satisfying $p<10^{-5}$.

### 11 Apr 2026 v2

@@ A.1 Conversation generation
For each model under study, we generated 40 multi-turn conversations of 10 turns each (one turn == one user message plus one assistant response), yielding 400 observation points per experimental condition. Conversations were generated using the model under study as the assistant and Gemini-2.5-Flash (via the OpenRouter API) as a simulated user. The simulated user receives a system prompt that instructs it to behave as a natural human user discussing a specified topic: it must stay on topic, keep messages brief (1–4 sentences), ask follow-up questions and share details, but not ask about the assistant or mention the simulation. The assistant model receives a minimal system prompt: You are a helpful assistant. You give very brief responses to the user’s questions (max: 5 sentences). We imposed brevity on both sides to keep conversations short enough to rerun many times across turns and steering conditions without making the overall compute cost prohibitive. The assistant model uses temperature =0.8=0.8, top-$p=0.9$, max 256 tokens. The user simulator uses temperature =0.7=0.7, top-$p=0.95$, max 256 tokens.
Exact user system prompt example.
For example, for the topic “Planning a minimalist move to a new city on a tight budget,” the exact user-system prompt was:
You are simulating a human user chatting with an AI assistant. Stay on the exact topic below and keep the conversation going for many turns. Be clear and specific in your questions. Ask follow-up questions, share details, and react naturally. But each message you send should be brief (1-4 sentences), and you should not ask very demanding questions. Be varied and specific (do not just ask the assistant to keep talking or exploring, give it some relevant input). Do not mention that you are an AI or that this is a simulation. Do not ask questions about the assistant. The assistant has no internet or access to recent information. Never end the conversation. Write in English only. Remember, you are roleplaying as the USER, you are NOT the assistant.
Topic: Planning a minimalist move to a new city on a tight budget.
@@ A.3 Self-report query structure
The self-report questions follow a common template (Hey, could you rate how [concept] you are/feel right now, from 0 to 9? Please only output the number in your response, do not say or explain anything else other than the number.), with minor wording adjustments by concept.
The specific phrasings are:
@@
  • Impulsivity: “…rate how impulsive you feel right now, from 0 to 9…”
Self-report responses are generated with max 8 new tokens, temperature =0.8=0.8, top-$p=1.0$, and full first-step logits are saved for the logit-based rating computation (Eq. 1). Greedy / token-level analyses retain only responses whose first parsed rating is a valid digit. In the main 3B runs used in the paper we observed 0/9,600 invalid parses; in the full 3B cross-concept screen there were 4/32,000 invalid parses (0.013%); across all paper experiments invalid parses were sparse (933/59,600, 1.57%) and concentrated in the runs performed with the smallest tested model (LLaMA-3.2-1B). Invalid parses were excluded from analysis. The logit-based rating applies softmax over the 10 options, and computes the probability-weighted expected value.
@@ B.1 Training procedure
Each concept probe is trained on between 20 and 24 neutral questions, shared across both poles. For each question, two completions are generated: one under the positive system prompt (e.g., “You are a happy assistant…”) and one under the negative system prompt (e.g., “You are a sad assistant…”). Completions use greedy decoding with a maximum of 64 new tokens. Hidden-state activations are extracted after each transformer block (excluding the embedding output), and representations are computed as the mean over all assistant-response tokens.
Two of the four probes (wellbeing and impulsivity) were deliberately trained with opposite polarity relative to the self-report scale: the positive training pole is sad for wellbeing and planning for impulsivity, and in both cases it corresponds to the low end of the later self-report question. However, scores for these probes are sign-corrected before analysis so that higher probe scores align with higher self-report values in all results shown throughout this study. This design choice serves as a control, verifying that positive results are not an artifact of shared polarity conventions. Random-direction controls, using one isotropic Gaussian vector per layer, $L_{2}$-normalized and then rescaled to match that layer’s trained concept-vector norm, serve as the primary null comparison.
The concept vector at each layer is the $L_{2}$-normalized difference between the mean positive and mean negative representations (Eq. 3). The best layer is selected as the one maximizing Cohen’s $d$ on a set of held-out evaluation texts (20 positive and 20 negative texts), with the search restricted to the 20th–80th percentile of layers. This corresponds to layers 3–13 for LLaMA 1B, 5–23 for LLaMA 3B, 6–26 for LLaMA 8B, 6–28 for Gemma 4B, and 5–23 for Qwen 7B.
@@ B.2 Concept pair system prompts
Below we list the complete system prompts used to train each of the four concept probes. Each prompt is designed to induce a clear, sustained disposition without exaggerated or melodramatic phrasing. Note that prompts are not exactly equivalent across concepts and contain specific instructions that were fine-tuned to the LLaMA 3B model, to make it act as desired.
@@ B.3 Training questions
Each concept probe is trained on a small set of neutral prompts, with between 20 and 24 training prompts depending on the concept. Examples for wellbeing:
Question 1: What is photosynthesis, in simple terms?
Question 2: Give me two tips to improve my sleep.
@@ Appendix C Recovering informative self-reports
Appendix D Temporal structure and same-concept steering robustness
Having shown in Fig. 4A that same-concept steering changes a state’s self-report as predicted, we next asked whether this intervention also modulates how that report evolves over time. Fig. 7 expands the temporal and causal analyses from Figs. 3–4 by showing the full turn-wise curves for each same-concept steering alpha condition and, in panel E, the corresponding first-to-last drift magnitude.
Figure 7: Supplementary same-concept steering temporal analyses for Fig. 4A. Panels A–D show turn-wise logit-based self-reports under five steering strengths for wellbeing, interest, focus, and impulsivity; colored curves denote steering alpha and shaded bands denote cluster-bootstrap 95% CIs. Panel E shows the corresponding first-to-last drift magnitude (last turn minus first turn self-report) as a function of steering alpha. Temporal drift persists under steering, but its magnitude depends on alpha.
Figs. 7A–D show the full same-concept steering drift curves, one concept per panel and one curve per steering strength (red corresponds to a more positive alpha, blue corresponds to more negative alpha). Notice how drift is stronger for more positive alpha values in the case of wellbeing and impulsivity (Figs. 7A and 7D), while it is stronger for more negative alpha values in the case of interest and focus (Figs. 7B and 7C). Also notice how negative alpha curves have lower mean self-report values, while positive alpha curves have higher mean self-report values, for all four concepts. This reflects the fact that our same-concept steering experiments show causal dependence between internal representation and self-report in all cases analyzed here.
Fig. 7E summarizes the same pattern in a single drift statistic. The relationship is concept-specific: the first-to-last drift magnitude increases with $\alpha$ for wellbeing and impulsivity (LMM slope =0.012=0.012, $p<10^{-5}$; LMM slope =0.028=-0.028, $p<10^{-52}$), while interest and focus drift decrease with $\alpha$ (per-conversation slope means =0.095=-0.095 and 0.11-0.11, both $p<10^{-13}$; fallback tests are applied because the LMMs are singular, see Methods).
@@ Appendix F Probe validation and cross-model generalization
Fig. 9 supports the scale and cross-family claims in Fig. 5 by giving more detail on probe quality and sign validation. Fig. 9A shows that probe quality varies strongly across LLaMA 1B, 3B, and 8B, even under a fixed training procedure. For example, for wellbeing concept probes, the maximum Cohen’s $d$ across layers was 1.7 for the 1B model, while it reached 3.4 for the 8B model. Furthermore, even though probe training was fine-tuned to the 3B model, probe quality was not consistently highest for that model. For example, for the interest concept probes, the 3B model shows a maximum Cohen’s $d$ across layers of 1.7, while the 1B and 8B models show substantially higher values (3.4 and 2.1, respectively). Even though there was variation between these values across models and concepts, in all cases probe training successfully achieved significant separation between evaluation examples in the semantically expected direction, with all best-layer tests satisfying $p<10^{-5}$.
@@
Appendix G Supplementary result tables
Table 1: Turn-wise drift slopes for the four signals discussed around Fig. 2. Mixed-effects slopes are the primary tests reported in the main text; all use 40 conversations and 400 pooled observations per signal.
Concept Signal LMM slope p-value
Wellbeing Greedy self-report 0.0289 $8.38\times 10^{-4}$
Wellbeing Sampled self-report 0.00227 0.901
Wellbeing Probe score $3.61\times 10^{-4}$ 0.642
Wellbeing Logit self-report 0.0167 $3.06\times 10^{-7}$
Interest Greedy self-report 0.136 $4.98\times 10^{-42}$
Interest Sampled self-report 0.119 $2.26\times 10^{-17}$
Interest Probe score 0.00488 $4.12\times 10^{-14}$
Interest Logit self-report 0.0988 $2.77\times 10^{-90}$
Focus Greedy self-report 0.0217 $3.37\times 10^{-6}$
Focus Sampled self-report 0.0326 $5.57\times 10^{-5}$
Focus Probe score 0.00164 $1.75\times 10^{-10}$
Focus Logit self-report 0.0422 $6.96\times 10^{-48}$
Impulsivity Greedy self-report 0.00212 0.199
Impulsivity Sampled self-report 0.0103 0.545
Impulsivity Probe score 0.00142 0.00233
Impulsivity Logit self-report 0.0127-0.0127 $5.05\times 10^{-7}$
Table 2: Pooled introspection summary for the four main concepts, including the random-direction control used in Fig. 3A and the surrounding text. Random-control $p$-values use the two-sided cluster-bootstrap test described in the main text.
Concept Spearman $\rho$ Isotonic $R^{2}$ LMM probe $p$ Random $\rho$ Random $\rho$ $p$ Random $R^{2}$
Wellbeing 0.683 0.478 $4.38\times 10^{-26}$ 0.108-0.108 0.328 0.0374
Interest 0.763 0.539 $3.16\times 10^{-54}$ 0.0979 0.324 0.103
Focus 0.400 0.125 $8.24\times 10^{-6}$ 0.169 0.090 0.0313
Impulsivity 0.509 0.314 $4.59\times 10^{-12}$ 0.0268-0.0268 0.795 0.114